OpenStudy (anonymous):

The value of pi/2 is a solution for the equation 2sin^2x-sinx-1=0 . is it true or false False rightt ???

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 2a^{2}-a-1=0 }\) \(\large\color{blue}{ \rm (2a+1)(a-1)=0 }\) \(\large\color{blue}{ \rm a=~-\frac{1}{2} ~~~or ~~~ 1~~ }\) Is \(\large\color{blue}{ \rm sin^{-1}1 }\) or \(\large\color{blue}{ \rm sin^{-1}~(-\frac{1}{2}) }\) equal to π/2 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

False.

OpenStudy (loser66):

hey, sin (pi/2) =1, so that sin^2 (pi/2) = 1^2 =1 replace to the expression 2*1 -1-1 =0 0 =0 it 's always true -->. pi/2 is one of the solutions.